### v1

# Large Language Models Can Learn Temporal Reasoning

Abstract
Large language models (LLMs) learn temporal concepts from the co-occurrence of related tokens in a sequence. Compared with conventional text generation, temporal reasoning, which reaches a conclusion based on mathematical, logical and commonsense knowledge, is more challenging. In this paper, we propose TempGraph-LLM, a new paradigm towards text-based temporal reasoning. To be specific, we first teach LLMs to translate the context into a temporal graph. A synthetic dataset, which is fully controllable and requires minimal supervision, is constructed for pre-training on this task. We prove in experiments that LLMs benefit from the pre-training on other tasks. On top of that, we guide LLMs to perform symbolic reasoning with the strategies of Chain of Thoughts (CoTs) bootstrapping and special data augmentation. We observe that CoTs with symbolic reasoning bring more consistent and reliable results than those using free text.
1 Introduction
As one of the fundamental abilities, temporal reasoning plays an important role in human perception. It is not just about understanding basic concepts such as order or duration; it extends to more intricate aspects, e.g., making sense of complex narratives or discovering causal relations. Recently, large language models (LLMs)  (Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023a) have shown remarkable performance in complex reasoning tasks  (Huang and Chang, 2023). However, they still can not perform temporal reasoning very well as we expected  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023). Compared with other singular types, temporal reasoning is a form of integrated task built upon foundations like mathematics, logical and commonsense knowledge  (Zhou et al., 2019; Qin et al., 2021).
Figure 1: Our TempGraph-LLM framework has two steps: text-to-tempgraph translation and neuro-symbolic reasoning.
Recent work mainly adopts the general methods to investigate and improve the temporal reasoning capability of LLMs. For example,  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023) benchmark leading LLMs on different temporal reasoning tasks with standard prompting, few-shot learning and Chain of Thoughts (CoTs)  (Wei et al., 2023a).  (Li et al., 2023; Yuan et al., 2023) introduce pre-defined Python programs/rule-based templates to perform supervised fine-tuning (SFT). Similarly,  (Wei et al., 2023b) design several special types of prompts for in-context learning (ICL). In addition,  (Tan et al., 2023a, b) adopt some extra strategies, which include span extraction pre-training, instruction tuning and reinforcement learning.
Despite the effectiveness of such methods, they ignore the intrinsic nature of temporal reasoning, and cannot guarantee the correctness of intermediate steps  Saparov and He (2023); Yang et al. (2023b). Human performs temporal reasoning on a timeline of events which are aligned with the entities and relations. These temporal concepts (e.g., ordering, duration, frequency, typical time) are then rigorously defined based on the timeline information. Thus, a natural solution for LLMs to reach human-level performance is to first extract a timeline from the text, and then to perform deliberate reasoning on it. However, the high-quality extraction of aligned timeline is a challenging task itself due to lack of ground truth. To solve this problem, we propose a synthetic dataset, which is fully controllable and requires minimal supervision. We show in experiments that the capability of timeline extraction learned on our dataset can be transferred to other tasks.
Given the reliable timeline, we still need to solve two problems to guide LLMs to perform rigorous reasoning on that. (1) How to introduce the necessary arithmetic and commonsense knowledge involved in temporal reasoning? (2) How to guide LLMs to perform rigorous reasoning? Inspired by retrieval-augmented generation (RAG)  (Lewis et al., 2021), we first extract all the time expressions, and then generate all the necessary information (e.g., the temporal relation and time gap of the timestamps, the relative order of the time gaps). As for (2), there are always two roadmaps: (i) translating natural language into logical statements and using symbolic inference engine for reasoning  Pan et al. (2023); (ii) directly using LLMs as the reasoning engine  (Zhu et al., 2023). For (i), the difficulty lies in accurate translation  Yang et al. (2023c) and the limited expressive power of formal logic. For (ii), there is no guarantee for the generated text especially when training data are insufficient  (Yang et al., 2023b). To solve these problems, we propose a symbolic integration bootstrapping method to generate reliable intermediate steps for supervised fine-tuning. Furthermore, we design two types of data augmentation strategies to mitigate data deficiency for neuro-symbolic reasoning tasks.
To be specific, our contributions are summarized as follows:
  • We propose a new paradigm, TempGraph-LLM, for text-based temporal reasoning. In this framework, we first translate the context into a temporal graph, and then perform symbolic reasoning on it. Extensive experiments prove that it brings better performance than convetional methods.
  • We design two approaches including chain of thoughts bootstrapping and data augmentation to guide LLMs to utilize symbolic reasoning, which brings better performance than free text.
  • We present a pipeline to create a synthetic dataet for question answering that requires temporal reasoning. It is fully controllable and requires minimal supervision for context-tempgraph alignment. We show in experiments that pre-training on our dataset benefits LLMs on other tasks.
2 Dataset Construction
We propose a pipeline for the construction of our synthetic temporal QA dataset called TempGraph QA. Each sample, as shown in Table 1, is in the form of (story, tempgraph, question, answer). The tempgraph is a subgraph extracted from a large temporal knowledge graph, YAGO11k (Dasgupta et al., 2018). The number of events in each tempgraph is restricted into a certain range. By using some rule-based templates, we obtain reliable question and answer pairs from the graph. Finally, each story is translated from the tempgraph by GPT-3.5. Below we show the details for all steps.
TempGraph (sub; rel; obj; start/end; time):
[…]
[2] (John Thompson owned Pearl Network) starts at 1942;
[3] (Sophia Parker was married to John Thompson) starts
at 1947; [4] (John Thompson was married to Sophia Parker)
starts at 1947; [5] (Sophia Parker was married to John
Thompson) ends at 1953; [6] (John Thompson was married
to Sophia Parker) ends at 1953; [7] (John Thompson owned
Pearl Network) ends at 1967; [8] (John Thompson died in
Riverside) starts at 1988; [9] (Sophia Parker died in
Lancaster) starts at 1995.
Graph-based Story (from GPT-3.5):
[…]
boy named John Thompson was brought into the world
in the year 1921. Growing up, he had a vibrant spirit
and an adventurous soul. $\cdots$
Graph-based QA (from Rule-based Python Script):
[…]
Pearl Network) or (John Thompson was married to
Sophia Parker)?
A: (John Thompson owned Pearl Network).
Q: True or false: event (John Thompson owned Pearl
Network) was longer in duration than event (Sophia
Parker was married to John Thompson)?
A: True.
$\cdots$
Table 1: Each sample from TempGraph QA is in the form of (story, tempgraph, question, answer).
Reasoning Type Question Answer
Sequencing Which event occurred first, {event A} or {event B}? {event A}/{event B}
Given the following five events: {event A}, {event B}, {event C}, {event D}, {event A}/{event B}/
{event E}, which event is the first/second/third/fourth/fifth one in chronolog- {event C}/{event D}/
ical order? {event E}
Duration How long did the {event A} last? {duration of event A}
Temporal How much time passed between the start of {event A} and the {gap between event A
Relation start of {event B}? and B’s start time}
What happened right before/after {event A} started? {event B} right before/
after {event A}
Facts When did the {event A} occur? {event A’s start time}
Extraction
Simultaneity True or false: {event A} and {event B} happened at the same year? True / False
True or false: {event A} was still happening when {event B} started? True / False
Comparative True or false: {event A} was longer in duration than {event B}? True / False
Analysis
Table 2: Reasoning types with corresponding questions and answers in TempGraph QA.
Step 1: Graph Splitting & Masking   Existing temporal knowledge graph datasets (Leetaru and Schrodt, 2013; Dasgupta et al., 2018; García-Durán et al., 2018) have tens of thousands of events, making it impossible to convert them all at once. Given a specific entity, we extract its third-order neighbors and all the events happening between them. To avoid the discrepancy caused by the number of events, we set a certain range and discard the non-compliant subgraphs. Since we want LLMs to do reasoning instead of memorization, it is ensured that no overlapping exists between the facts in training and test samples. We finally obtain 500 tempgraphs for training and another 100 for test. In addition, we consider the noticeable leakage problem of LLMs, where the prior knowledge of test data has been implicitly obtained from the extensive pre-training. Thus, a masking strategy, i.e., changing entities names into random ones of the same type, is adopted. We generate a global mapping for entity names with GPT-3.5. To avoid confusion, we only choose obscure names that do not exist in the original large knowledge graph.
Step 2: Graph-based Open QA Creation   Based on the given tempgraph, we generate a story and multiple QAs. We first ask GPT-3.5 to write a story based on the graph with the requirements to include all the facts. It is observed that GPT-3.5 tends to ignore the end time of some events in the created story. To solve this problem, we split the starting and ending of a persistent event into two transient ones. To obtain a comprehensive benchmark, we consider all types of temporal reasoning, which include sequencing, duration, frequency, simultaneity, temporal relation, comparative analysis and facts extraction. Given these categories, we design nine different question types (shown in Table 2), and write a rule-based Python script to determine the answers from the tempgraph.
Step 3: Quality Control   Although LLMs are good at story generation, there is a risk of misalignment between the given tempgraph and the generation text. To avoid the interference caused by misalignment, we propose an automatic way for verification. Given a generated story, we ask LLMs the time of each event in the graph. If LLMs cannot give the correct answer, we remove the possibly missing events in the graph. Since this strategy tends to simplify the reasoning, we manually verify the existence of the possibly missing events for both training and test sets. Note that only the quality of story-tempgraph alignment needs control, since all the QAs are generated from the tempgraph with specific rules. We show dataset statistcis in Appendix A, and all the prompts involved in Appendix C.
3 TempGraph-LLM
Motivated by human perception, we propose a new paradigm called TempGraph-LLM. We first translate the text into a temporal graph, and then guide LLMs to perform symbolic reasoning on it.
3.1 Text-to-TempGraph Translation
TempGraph (TG) serves as the foundation of the following symbolic reasoning process. Although LLMs can extract them from the given context without fine-tuning, we observe a misalignment between the generated graph and pre-defined QAs. Thus, we provide a deliberate pipeline for high-quality TG translation.
3.1.1 Pre-training
In the TempGraph QA, each story is corresponding to a verified TG. Similarly, for other data, where the ground truth of TG is avaiable, we can pre-train LLMs on them. Since language models (LMs) can only deal with sequence of tokens, we sort all the events in the graph according to their time. In addition, to deal with both transient and persistent events, we consider the starting and ending of a persistent event as two transient events. With these steps, we transfer the TG into a timeline with the alignment between entities, relations and times. We perform supervised fine-tuning (SFT) using Llama-2 model  (Touvron et al., 2023b) with Low-Rank Adaptation (LoRA)  (Hu et al., 2021). The input and output of the model are the story and aligned timeline, respectively. In experiments, we observe benefits from the pre-training on our data for other tasks.
3.1.2 Fine-tuning
Generally, the biggest challenge for fine-tuning on real-world datasets is lack of ground truth. Instead of directly using the estimation from LLMs, we provide a pipeline for better alignment.
Entity & Relation Extraction   To align the entities and relations in generated TGs and pre-defined QAs, we use a prompt with requirements on including the ones in pre-defined QAs when extracting from the context. To obtain the entities and relations in pre-defined QAs, we consider two ways: parsing the QAs with rules or using LLMs. The former is more efficient and reliable when the QAs are generated by templates  (Chen et al., 2021; Tan et al., 2023a), while the latter should be used when the QAs are in free text  (Zhou et al., 2019; Ning et al., 2020).
Temporal Info Identification   Compared with conventional reasoning tasks, temporal information is one of the most important and unique components here. To identify all the time expressions, LLMs are first used to extract them in the context, and then a rule-based Python script  (Yang et al., 2023a) is adopted for time-normalization and invalid output filtering. For example, we convert "Dec. 5th, 2004" into "2004-12-05", "July 1989" into "1989-07-XX", and "1973" into "1973-XX-XX", respectively. For the valid normalized time expressions, we can easily obtain their temporal relations and time gaps, e.g., "1974 is before 1983", "the time gap between May 2004 and Oct 2008 is 4 years and 5 months", which are introduced in temporal reasoning as external knowledge.
Graph Construction & Verification   We generate the graph using LLMs with in-context learning (ICL). To be specific, we provide in the prompt several examples which include a story and extracted entities, relations and time expressions as input and the corresponding TG as output. The TG should only contain the extracted elements to guarantee the alignment with pre-defined QAs. We classify the stories according to whether they contain explicit expressions such as "July 1989", "2003–2005", "the late 1900s". For those without explicit expressions, we ask LLMs to provide the temporal order of different events. We further automatically verify the generated TGs by asking LLMs to answer the pre-defined QAs based on the graph. For those graphs where LLMs give the wrong answers, we manually check and correct them.
3.2 Neuro-symbolic Temporal Reasoning
Given the reliable TGs, we train LLMs to perform symbolic reasoning with SFT enhanced by Chain of Thoughts (CoTs) bootstrapping and data augmentation.
3.2.1 Bootstrapping Chain of Thoughts
It is observed that SFTs on reliable CoTs bring better performance in reasoning tasks than those on standard Input/Output prompts  (Wang et al., 2023; Ho et al., 2023). However, it is challenging to obtain reliable CoTs from LLMs for symbolic reasoning tasks  (Saparov and He, 2023). Inspired by  (Zhu et al., 2023), we provide several manually verified examples to LLMs as ICL, and ask LLMs to generate CoTs for the new samples. We bootstrap the generations and only accept the ones with correct final answers.
To be specific, we first classify the pre-defined QAs into several categories $\{Q_{1},Q_{2},\cdots,Q_{m}\}$ with LLMs. For each category $Q_{i}$, we randomly choose $N_{i}$ samples $\{(g_{j},e_{j},q_{j},a_{j})\}^{N_{i}}_{j=1}$ and ask LLMs to provide symbolic CoTs $\{c_{j}\}$ given the TG $\{g_{j}\}$, external knowledge $\{e_{j}\}$ and question $\{q_{j}\}$. These CoTs will be manually verified and used as few-shot ICL examples. When given new samples $(g_{j^{\prime}},e_{j^{\prime}},q_{j^{\prime}},a_{j^{\prime}})$, we bootstrap CoTs with final answers $(c^{k}_{j^{\prime}},\hat{a}^{k}_{j^{\prime}})$ from LLMs, and only accept the one leading to correct answers, i.e., $\hat{a}^{k}_{j^{\prime}}=a^{*}_{j^{\prime}}$. Compared with generating CoTs with templates  (Saparov and He, 2023; Yuan et al., 2023), our method is more flexible and robust. In addition, inspired by  (Wang et al., 2023), we compare different type of CoTs using contrastive learning. The core idea is to choose CoTs that make the correct answer more plausible than other candidates.
$c^{*}=\operatorname{argmax}\log P\left(a^{*}\mid g,e,q,c\right)+\gamma G\left(c\right)$ (1)
$G\left(c\right)=\log\frac{P\left(a^{*}\mid g,e,q,c\right)}{\bar{P}_{\{a^{\prime}\in A^{\prime}\}}\left(a^{\prime}\mid g,e,q,c\right)}$ (2)
$P\left(a\mid g,e,q,c\right)=\frac{1}{|a|}\underset{t_{l}\in a}{\sum}P\left(t_{l}\mid g,e,q,c,t_{<l}\right)$ (3)
where $g,e,q,c$ denote the TG, external knowledge, question, and CoTs, respectively; $a^{*}$ denotes the correct answer; $a^{\prime}$ denotes the wrong answer from a candidate set $A^{\prime}$; weight $\gamma$ is a hyperparameter; $G(\cdot)$ denotes the contrastive plausibility growth; conditional probability $P(\cdot)$ is calculated in a way similar to perplexity, $t_{l}$ denotes the $l$-th token in the answer $a$, and $t_{<l}$ denotes the sequence of tokens before $t_{l}$.
3.2.2 Data Augmentation
Compared with conventional tasks, reasoning suffers more from data insufficiency since more information, such as evidence, arguments, and logics, are involved in intermediate steps  (Huang and Chang, 2023). To address the issue, we propose two special strategies for data augmentation.
Learning with Disturbances   Essentially, our framework involves two steps: text-to-tempgraph translation and symbolic temporal reasoning. At the test stage, reasoning happens on the noisy graphs, which can impact the overall performance. Specifically, for the reasoning module, we train with ground truth/verified graphs while testing on the estimations. To improve the robustness of reasoning, we introduce some disturbances on the graphs during training. Note that the type of disturbances should be carefully designed in order to avoid confusing the model. We investigate two types of disturbances: remove irrelevant edges or replace edges by changing relation names, e.g., "owned" $\rightarrow$ "run", "was affiliated to" $\rightarrow$ "belonged to", "played for" $\rightarrow$ "joined". An edge is considered as irrelevant if it is not involved in both QA and CoTs.
$F_{irr}\in\{F:P(c|g_{\setminus\{F\}},e,q)\approx P_{0}\}_{F\in g}$ (4)
$F^{\prime}\in\{f_{R}(F):P(c|f_{R}(g),e,q)\approx P_{0}\}_{F\in g}$ (5)
TempGraph (Adding Disturbances):
[…]
[2] (John Thompson run Pearl Network) starts at 1942;
[3] (Sophia Parker and John Thompson became life
partner) starts at 1947; [4] (John Thompson and Sophia
Parker became life partner) starts at 1947; [5] (Sophia
Parker and John Thompson became life partner)
ends at 1953; [6] (John Thompson and Sophia Parker)
became life partner) ends at 1953; [7] (John Thompson
run Pearl Network) ends at 1967.
Graph-based QA (No Change):
[…]
Network) was longer in duration than event (Sophia
Parker was married to John Thompson)?
CoT (symbolic): The duration for each event can be
calculated as follows:
(John Thompson owned Pearl Network) starts at
1942, ends at 1967, 1967 - 1942 = 25
(Sophia Parker was married to John Thompson) starts
at 1947, ends at 1953, 1953 - 1947 = 6
25 is greater than 6 , thus, the answer is True.
A: True.
TempGraph (Inductive Transformation):
[…]
[2] (James Brown owned Iris Inn) starts at 1952;
[3] (Ella Perry was married to James Brown) starts
at 1957; [4] (James Brown was married to Ella Perry)
starts at 1957; [5] (Ella Perry was married to James
Brown) ends at 1963; [6] (James Brown was married
to Ella Perry) ends at 1963; [7] (James Brown owned
Iris Inn) ends at 1977; [8] (James Brown died in
Auckland) starts at 1998; [9] (Ella Perry died in
Monaco) starts at 2005.
Graph-based QA (Inductive Transformation):
[…]
was longer in duration than event (Ella Perry was
married to James Brown)?
CoT (symbolic): The duration for each event can be
calculated as follows:
(James Brown owned Iris Inn) starts at 1952, ends at
1977, 1977 - 1952 = 25
(Ella Perry was married to James Brown) starts at
1957 , ends at 1963, 1963 - 1957 = 6
25 is greater than 6 , thus, the answer is True.
A: True.
Table 3: We propose two special data augmentation strategies (adding disturbances and inductive transformation) for the neuro-symbolic reasoning.
where $P_{0}:=P(c|g,e,q)$ denotes the original conditional probability, irrelevant event $F_{irr}$ will be randomly removed from $g$, and disturbed event $F^{\prime}$ is obtained from the global mapping of relation name changing $f_{R}(\cdot)$.
Inductive Transformation  In symbolic reasoning, a rule is considered as inductive if it is generalizable to different entities and times  (Xiong et al., 2022, 2023). For LLMs, this might be difficult to understand without enough training data since their fundamental principle is to learn the co-occurrence of tokens. To solve this problem, we introduce a new approach called inductive transformation. We globally map all the entity names in training data to some random names of the same type (e.g., "Knox Cunningham" $\rightarrow$ "Olivia Davis"; "Sydney" $\rightarrow$ "Cardiff"), and change the times based on a global time offset (e.g., the offset is "6 years": "June 1994" $\rightarrow$ "June 2000"; "2008" $\rightarrow$ "2014"). For each relation, we generate these random names with LLMs by providing several examples of existing entity names. To avoid confusing the model, we only use obscure names that do not exist in our dataset.
$g^{\prime\prime}=f_{T}(f_{E}(g)),\ e^{\prime\prime}=f_{T}(e)$ (6)
$q^{\prime\prime}=f_{T}(f_{E}(q)),\ c^{\prime\prime}=f_{T}(f_{E}(c))$
$a^{\prime\prime}=f_{T}(f_{E}(a))$
where $f_{E}(\cdot),f_{T}(\cdot)$ denote the global mapping of entity name changing and time changing, respectively.
4 Experiments
We aim to answer the following research questions in our experiments: (1) Can we guide LLMs to perform more reliable symbolic reasoning with the strategies of bootstrapping CoTs and data augmentation? (2) Can our two-step framework lead to better temporal reasoning performance? (3) Can LLMs benefit from our data on other reasoning tasks?
4.1 Experimental Setup
In addition to TempGraph QA, we consider the existing dataset, TimeQA  (Chen et al., 2021), which is based on the texts from Wikipedia. It contains two difficulty levels: The easy-level split tends to be information extraction while the hard-level split involves temporal reasoning, i.e., understanding the relation between different temporal expressions. We evaluate our framework on both datasets with the metrics of token-level F1, exact match (EM) and perplexity-based accuracy (Acc). Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider the accuracy based on perplexity for LLM evaluations. Compared with other LMs, LLMs are pre-trained on a wide variety of tasks, making them more generalizable at the cost of less conforming to the required format. This property actually hurts their performance under the commonly used metrics such as EM and F1. To solve this problem, we select the answers from a candidate set based on their perplexity from the LLM. For questions with multiple correct answers, we follow the strategy proposed in  (Chen et al., 2021) to get the best F1, EM and Acc among them. For TempGraph QA, we first calculate the metrics for each category, and then use the average as final scores to mitigate question category imbalance.
We compare our framework with leading LLMs including Llama2, GPT-3.5 and GPT-4. Llama2 Touvron et al. (2023b) is one of the most accepted open-source LLMs, while GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI, 2023) are two powerful closed-source models. We run Llama2 series on local machines (4-bit quantization is adopted for 70B version due to computational resource limitation), and use APIs provided by OpenAI for GPT series. We evaluate their few-shot ICL performance on the test set. To avoid LLMs generating free text which affects the evaluation with these metrics  (Yang et al., 2023a), we provide candidate answers for all questions under the ICL setting. The model versions and prompt templates are shown in Appendix B and C, respectively.
4.2 Implementation Details
We use Llama2-13B as the baseline due to computation resource limit. In experiments, we proved that our framework has comparable or even better temporal reasoning performance than one of the most powerful LLMs, GPT-4. We inject two adapters with selectors into the Llama2 model for the text-to-tempgraph translation and symbolic temporal reasoning. The adapters are trained in parallel. For inference, we first translate the original story into a temporal graph, and then perform symbolic temporal reasoning on it, i.e., the adapters are used in sequence. In addition, we use GPT-3.5 for the story creation & verification in TempGraph QA, entity and relation extraction, temporal info identification, graph construction & verification on other datasets, and symbolic CoTs bootstrapping. These examples used in ICL are created by GPT-4 with manual verification. We choose these models due to the high quality of their generation. All the prompt templates are shown in Appendix C.
4.3 Main Results
Can our strategies guide LLMs to perform more reliable symbolic reasoning?   Figure 2 shows the comparison between ICL with CoTs, SFT with CoTs, bootstrapping CoTs, and data augmentation on TempGraph QA. To obtain a fair comparison, all the reasoning happens on the estimated graphs obtained from our text-to-graph translation module. It can be seen that LLMs learn better temporal reasoning with SFT than ICL. By providing CoTs with symbolic reasoning, the model performance gets further enhanced. There is an additional performance boost when using our data augmentation strategies. Furthermore, inspired by  (Wang et al., 2023), we manually check 100 CoTs generated by different strategies. Evaluators are asked to classify the errors into four types (using wrong info, logical inconsistency, external knowledge error, and temporal graph error). We measure the level of agreement among the three evaluators, and show the error rates from different types in Table 4. It can be seen that LLMs with ICL easily use either wrong information or erroneous external knowledge in reasoning. This issue can be alleviated with SFT, and we shows that using symbolic reasoning is better than free-text CoTs. Our data augmentation strategies further reduce all types of errors, and provides a promising direction for LLMs on reasoning tasks.
Figure 2: Performance comparison between different strategies of CoT generation on TempGraph QA.
Strategy ER-T1 ER-T2 ER-T3 ER-T4
ICL 0.13 0.13 0.31 0.10
SFT 0.04 0.17 0.13 0.10
SFT + bs 0.06 0.13 0.03 0.08
SFT + bs + aug 0.03 0.05 0.04 0.05
Table 4: Human evaluation on the CoTs generated by different strategies. T1: using wrong info; T2: logical inconsistency; T3: external knowledge error; T4: temporal graph error.
Model TempGraph QA TimeQA
Easy-mode Hard-mode
EM F1 Acc EM F1 Acc EM F1 Acc
GPT-3.5 (ICL-SP) 0.559 0.699 - 0.640 0.668 - 0.512 0.506 -
GPT-3.5 (ICL-CoT) 0.698 0.788 - 0.565 0.603 - 0.436 0.464 -
GPT-4 (ICL-SP) 0.704 0.818 - 0.716 0.742 - 0.571 0.546 -
$*{}^{*}$ (ICL-CoT) 0.798 0.833 - 0.662 0.693 - 0.618 0.636 -
Llama2-7B (ICL-SP) 0.402 0.596 0.447 0.352 0.408 0.367 0.341 0.404 0.354
Llama2-7B (ICL-CoT) 0.548 0.686 0.578 0.367 0.425 0.393 0.302 0.354 0.360
Llama2-13B (ICL-SP) 0.423 0.609 0.526 0.439 0.493 0.450 0.427 0.481 0.437
Llama2-13B (ICL-CoT) 0.626 0.762 0.668 0.518 0.572 0.535 0.434 0.490 0.480
Llama2-70B (ICL-SP) 0.578 0.736 0.665 0.583 0.627 0.631 0.493 0.537 0.551
Llama2-70B (ICL-CoT) 0.760 0.838 0.796 0.552 0.612 0.623 0.447 0.501 0.512
Llama2-13B (SFT-SP) 0.550 0.720 0.652 0.412 0.449 0.455 0.362 0.404 0.412
Llama2-13B (SFT-CoT) 0.646 0.722 0.705 0.550 0.586 0.564 0.332 0.391 0.379
Llama2-13B (TempGraph) 0.797 0.850 0.819 0.664 0.691 0.673 0.631 0.664 0.649
Table 5: Main results using different models and strategies. We report exact match (EM), token-level F1 scores, and perplexity-based accuracy (Acc). Results with * are evaluated on 1000 random test samples.
Figure 3: Performance comparison between different strategies on TimeQA. To obtain a fair comparison, we use Llama2-13B as the base model for all strategies.
Can our two-step framework lead to better temporal reasoning performance?   Table 5 shows the comparison between different models on both TempGraph QA and TimeQA. First, we observe that among all the LLMs with ICL, GPT-4 has the best temporal reasoning capability as expected. For Llama2 series, larger model has better performance due to advanced context understanding and improved generalization. We also find that CoTs not always bring better temporal reasoning. In both easy-mode and hard mode of TimeQA, CoTs hurt larger LMs such as GPT-3.5, GPT-4 and Llama2-70B due to unreliable intermediate results and hallucinations. More importantly, our model, which is based on Llama2-13B, shows a comparable or even better temporal reasoning performance than GPT-4 on both datasets. To prove the effectiveness of our framework, we also provide the results of some alternative strategies (ICL, SFT-SP, SFT-CoT), where SFT-SP and SFT-CoT denote supervised fine-tuning on standard Input/Output prompts and CoTs, respectively. Compared with our framework, there are no text-to-graph translation and symbolic reasoning involved in SFT-SP and SFT-CoT. It can be seen that the two-step framework brings a substantial improvement in model performance on different datasets. Compared with original text, we actually provide an easier path from temporal reasoning questions to their answers.
Can LLMs benefit from our data on other reasoning tasks? Figure 3 shows the performance comparison between different strategies on TimeQA (ICL-SP, ICL-CoT, pre-training on TempGraph QA, SFT on TimeQA). It can be seen that the performance of in-context learning (either with standard I/O prompt or CoTs) is much lower than pre-training and SFT. In our framework, LLMs can learn the capability of text-to-graph translation and symbolic reasoning, which brings better temporal reasoning. More importantly, we observe that per-training on synthetic data improves the model performance compared with ICL, which proves the rationality of design and construction of our dataset. Also, it can be concluded that these necessary capabilities in temporal reasoning are generalizable to different data distributions. Since the synthetic dataset is fully controllable and requires minimal supervision, we actually provide a way of improving the reasoning capability of LLMs.
4.4 Ablation Study
We ablate different modules to see their contributions to the model performance. Table 6 shows the performance comparison between different configurations where TG, EK, SR denote temporal graph, external knowledge and symbolic reasoning, respectively. To obtain a fair comparison, we use Llama2-13B as the base model for all configurations. We use supervised fine-tuning on CoTs (SFT-CoT) as baseline. We improve the model performance 2.9% in EM, 3.2% in F1 and 0.6% in Acc by adding TG. Then we show that bootstrapping CoTs with symbolic reasoning brings further enhancement of 4.8% in EM, 4.3% in F1 and 2.5% in Acc. These is a significant boost in EM (5.9%), F1 (4.1%) and Acc (5.9%), when using our data augmentation strategies. Finally, the introduction of external knowledge exhibits meaningful results: it increases EM by 3%/1.5%, F1 by 0.6%/1.2%, and Acc by 4%/2.4% without/with data augmentation. From the ablation study, we give some insights for LLMs: (1) LLMs can benefit from (temporal) graph which is intuitive, concise and structured. (2) Given a reliable graph, symbolic reasoning brings better performance in reasoning than free-text CoTs. (3) Data augmentation is necessary for LLMs to perform complex tasks such as temporal reasoning. (4) The introduction of external knowledge such as mathematics and commonsense can augment the generation of LLMs.
Config EM F1 Acc
Baseline (SFT-CoT) 0.646 0.722 0.705
TG 0.675 0.754 0.711
TG + SR(bs) 0.723 0.797 0.736
TG + SR(bs + aug) 0.782 0.838 0.795
TG + EK + SR(bs) 0.753 0.803 0.776
TG + EK + SR(bs + aug) 0.797 0.850 0.819
Table 6: Ablation Study Results on TempGraph QA. To obtain a fair comparison, we use Llama2-13B as the base model for all configurations.
5 Related Work
Text-based Temporal Reasoning   Recently, text-based temporal reasoning has gained substantial research focus  (Nylund et al., 2023; Hu et al., 2023; Gurnee and Tegmark, 2023; Liu et al., 2023). The vision here is to help LMs understand temporal concepts, which are fundamental elements in human knowledge. Existing methods mainly solve this problem with time-aware modeling. For example,  (Rosin et al., 2022; Pereira, 2022; Tan et al., 2023a) propose pre-training/fine-tuning strategies for robust temporal reasoning.  (Ning et al., 2019; Zhou et al., 2020; Yang et al., 2020; Zhou et al., 2021; Yang et al., 2023a) design auxiliary objectives to introduce external knowledge or extra temporal constraints. Although these methods make some progress, the underlying structure and logic of temporal reasoning are either ignored or not explicitly involved. Since LMs perform pre-training on human knowledge, the manual perspective of modelling might effectively improve their performance on this task.
Reasoning towards LMs   Although LLMs exhibit some emergent behaviors, it is still unknown whether they can actually perform reasoning and their capability of reasoning (Valmeekam et al., 2022; Wei et al., 2022; Huang and Chang, 2023). Existing methods that try to elicit or improve reasoning can be divided into two directions: reasoning-involved modelling or hybrid methods. For example,  (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023) use prompt engineering where some examples including intermediate reasoning steps are provided in context.  (Talmor et al., 2020; Nye et al., 2021; Wang et al., 2023; Ho et al., 2023) fine-tune LMs with these information before producing the final answers. On the other hand,  (Li et al., 2023; Kynoch et al., 2023; Zhuang et al., 2023; Xu et al., 2023) combine LMs with domain-specific tools, empowering the model to perform more complex tasks that requires reasoning. To the best of our knowledge, the exploration of the full capabilities of LLMs on temporal reasoning is still missing. Compared with other reasoning types, it is a form of integrated task that requires the knowledge of mathematics, logic and commonsense.
Neuro-symbolic Reasoning   Recent research shows advantages of combining neural networks with symbolic reasoning  (Sarker et al., 2021; Susskind et al., 2021; Hamilton et al., 2022). Knowledge graph is an important representation as the foundation of symbolic reasoning. Existing works on symbolic reasoning over KGs include  (Yang et al., 2017; Payani and Fekri, 2019a, b; Yang and Song, 2020; Qu et al., 2021). On top of that,  (Liu et al., 2022; Xiong et al., 2022, 2023) develop approaches for symbolic reasoning over temporal KGs. In this paper, we build the connection between text-based and temporal KG-based temporal symbolic reasoning. This connection brings the potential to extend these KG-based methods on text-based tasks. Some examples include  (Luo et al., 2023; Zhang et al., 2023; Yuan et al., 2023). Different from these methods, which limit their application to certain datasets, our framework has better generalization due to the text-to-graph translation part.
6 Conclusion
TempGraph-LLM, a new paradigm for language models, has been proposed to improve their performance on temporal reasoning. Besides producing the final answers, our framework also provides the temporal graph and symbolic reasoning process. Experiments indicate that TempGraph-LLM achieves better performance than existing methods. An interesting direction for future work is to extend it to more complex applications such as induction and abductive reasoning. Due to the graph structure and capability of symbolic reasoning, it is promising to improve the performance of LLMs on these tasks with good problem formalization and methodology extension.
Appendix A Dataset Statistics
In this section, we show statistics for our TempGraph QA dataset. Figure 4 shows the number of facts in training and test set, which determine the complexity of task to some extent. Figure 5 shows the types of questions in training and test set. Different types of questions are exemplified in Table 1. Among all the types, question 1 has the largest portion since it has five variants. Similarly, question 7 with two variants also has a larger portion. To mitigate question category imbalance, we first calculate the metrics for each category, and then use the average as final scores.
Figure 4: Number of facts in TempGraph QA.
Figure 5: Types of questions in TempGraph QA.
Appendix B Model Versions
All the model versions used in our experiments are listed below:
  • Llama2-7B (meta-llama/Llama-2-7b-hf)
  • Llama2-13B (meta-llama/Llama-2-13b-hf)
  • Llama2-70B (meta-llama/Llama-2-70b-hf)
  • GPT-3.5 (gpt-3.5-turbo)
  • GPT-4 (gpt-4-1106-preview)
Appendix C Example Prompts
In this section, we show example prompts used in our framework. To be specific, Table 7 shows an example of the graph-based story generation in TempGraph QA; Table 8 presents an example of the automatic story-tempgraph alignment verification in TempGraph QA; Table 9 provides an example of the temporal info identification in TimeQA; Table 10 illustrates an example of the graph construction in TimeQA; Table 11 gives an example of the automatic tempgraph-QA alignment verification in TimeQA; Table 12 depicts an illustrative case of the CoTs bootstrapping in TempGraph QA.
(Molly Adams was born in Seattle) starts at 1896
(Liam Thomas Dawson was born in Seattle) starts at 1896
(Molly Adams was married to Liam Thomas Dawson) starts at 1920
(Liam Thomas Dawson was married to Molly Adams) starts at 1920
(Liam Thomas Dawson won prize James Parker Prize for Advanced Biomedical
Studies) starts at 1946
(Liam Thomas Dawson won prize Champion Award in Biology or Science) starts at 1947
(Liam Thomas Dawson won prize Oakley Smith Prize) starts at 1948
(Liam Thomas Dawson won prize Member of the National Academy) starts at 1950
(Molly Adams was married to Liam Thomas Dawson) ends at 1957
(Liam Thomas Dawson was married to Molly Adams) ends at 1957
(Liam Thomas Dawson died in Kingsville, Texas) starts at 1984
Based on the above facts, generate a short story for me.
Table 7: Example of the graph-based story generation in TempGraph QA.
Once upon a time, in the vibrant city of Seattle, two remarkable individuals, Molly
Adams and Liam Thomas Dawson, were born in the year 1896. Little did they know that
their destinies were entwined from the very beginning.
Growing up in the same neighborhood, Molly and Liam developed a deep friendship
that blossomed into something more as they entered adulthood. In the year 1920, their
love story officially began as they exchanged vows and embarked on a journey of
companionship that would last for decades.
$\cdots$
When did the event (Molly Adams was married to Liam Thomas Dawson) end?
Table 8: Example of the automatic story-tempgraph alignment verification in TempGraph QA.
Knox Cunningham
Sir Samuel Knox Cunningham, 1st Baronet, QC (3 April 1909 – 29 July 1976) was a
Northern Irish barrister, businessman and politician. As an Ulster Unionist politician at
a time when the Unionists were part of the Conservative Party, he was also a significant
figure in United Kingdom politics as Parliamentary Private Secretary to Harold
Macmillan. His nephew was Sir Josias Cunningham.
$\cdots$
Extract all the time expressions such as ’June 1994’, ’1973’, ’late 1980s’.
Table 9: Example of the temporal info identification in TimeQA.
Knox Cunningham
Sir Samuel Knox Cunningham, 1st Baronet, QC (3 April 1909 – 29 July 1976) was a
Northern Irish barrister, businessman and politician. As an Ulster Unionist politician at
a time when the Unionists were part of the Conservative Party, he was also a significant
figure in United Kingdom politics as Parliamentary Private Secretary to Harold
Macmillan. His nephew was Sir Josias Cunningham.
$\cdots$
Construct a timeline for Knox Cunningham’s position. You should only consider these
time points (3 April 1909, 1930s, 1939, 1942, 1943, 1945, 1947, 1949, 1954, 29 July 1976).
Table 10: Example of the graph construction in TimeQA.
3 April 1909: Knox Cunningham was born.
1930s: He studied law and began his legal career.
1935: On 2 July 1935, he married Dorothy Enid Riley JP.
1939: Cunningham was called to the Bar by the Middle Temple.
1942: He was called to the Bar in Northern Ireland.
1943: Cunningham fought the Belfast West by-election.
1945: He contested the same seat in the general election.
1947: Knox Cunningham became involved with the World Alliance of YMCAs.
1949: He became Chairman of the National Council of the YMCA.
1954: Cunningham was elected to Orpington Urban District Council.
29 July 1976: Knox Cunningham passed away.
Which position did Knox Cunningham hold from May 1955 to Apr 1956?
Table 11: Example of the automatic tempgraph-QA alignment verification in TimeQA.
Example:
Timeline:
(Liam Mitchell was born in Harrison) starts at 1885
(Maddox Reynolds was born in Glenville) starts at 1893
(Liam Mitchell was married to Emma Scott) starts at 1916
(Emma Scott was married to Liam Mitchell) starts at 1916
(Liam Mitchell was married to Emma Scott) ends at 1918
(Emma Scott was married to Liam Mitchell) ends at 1918
(Maddox Reynolds was married to Liam Mitchell) starts at 1922
(Liam Mitchell was married to Maddox Reynolds) starts at 1922
(Maddox Reynolds was married to Liam Mitchell) ends at 1928
(Liam Mitchell was married to Maddox Reynolds) ends at 1928
(Liam Mitchell died in Boston) starts at 1941
(Emma Scott died in Oceanview) starts at 1984
Question: True or false: event (Liam Mitchell was married to Maddox Reynolds) was
longer in duration than event (Emma Scott was married to Liam Mitchell)?
Useful information:
1885 before 1893 before 1916 before 1918 before 1922 before 1928 before 1941
1918 - 1916 = 2
1928 - 1893 = 35
1928 - 1922 = 6
1941 - 1918 = 23
2 < 6 < 23 < 35
Answer: Let’s think step by step.
To determine whether the statement is true or false, we need to compare the durations
of (Liam Mitchell was married to Maddox Reynolds) and (Emma Scott was married to
Liam Mitchell) . The duration for each event can be calculated as follows:
(Liam Mitchell was married to Maddox Reynolds) :
(Liam Mitchell was married to Maddox Reynolds) starts at 1922
(Liam Mitchell was married to Maddox Reynolds) ends at 1928
1928 - 1922 = 6
(Emma Scott was married to Liam Mitchell) :
(Emma Scott was married to Liam Mitchell) starts at 1916
(Emma Scott was married to Liam Mitchell) ends at 1918
1918 - 1916 = 2
Now, we compare the duration of the two events: 6 is greater than 2 , thus, the answer is True .
Test:
Timeline:
(Chris Evans was born in Bristol, Connecticut) starts at 1948
(Oliver Jensen was born in Bristol, Connecticut) starts at 1948
(Chris Evans was married to Oliver Jensen) starts at 1970
(Oliver Jensen was married to Chris Evans) starts at 1970
(Chris Evans created Pearl Network) starts at 2005
(Chris Evans won prize Victory Achievement Award) starts at 2007
(Chris Evans was married to Oliver Jensen) ends at 2010
(Oliver Jensen was married to Chris Evans) ends at 2010
(Chris Evans created Pearl Network) ends at 2013
Question: True or false: event (Chris Evans owned Pearl Network) was longer in
duration than event (Oliver Jensen was married to Chris Evans)?
Useful information:
1948 before 1970 before 2005 before 2007 before 2010
1970 - 1948 = 22
2007 - 1948 = 59
2007 - 2005 = 2
2010 - 1970 = 40
2 < 22 < 40 < 59
Answer: Let’s think step by step.
Table 12: Example of the CoTs bootstrapping in TempGraph QA.