### v2

# Large Language Models Can Learn Temporal Reasoning

Abstract
While large language models (LLMs) have demonstrated remarkable reasoning capabilities, they are not without their flaws and inaccuracies. Recent studies have introduced various methods to mitigate these limitations. Temporal reasoning (TR), in particular, presents a significant challenge for LLMs due to its reliance on diverse temporal expressions and intricate contextual details. In this paper, we propose TG-LLM, a new framework towards language-based TR. To be specific, we first teach LLM to translate the context into a temporal graph (TG). A synthetic dataset, which is fully controllable and requires minimal supervision, is constructed for fine-tuning on this graph translation task. We confirm in experiments that the capability of TG extraction learned on our dataset can be transferred to other TR tasks and benchmarks. On top of that, we guide LLM to perform symbolic reasoning over the TG via Chain of Thoughts (CoTs) bootstrapping and special data augmentation strategies. We observe that CoTs with symbolic reasoning bring more consistent and reliable results than those using free-form text.
1 Introduction
As one of the fundamental abilities, temporal reasoning (TR) plays an important role in human perception. It is not just about understanding basic concepts such as order or duration; it extends to more intricate aspects, e.g., making sense of complex narratives or discovering causal relations. Recently, large language models (LLMs)  (Ouyang et al., 2022; OpenAI et al., 2023; Touvron et al., 2023b) have shown remarkable performance in complex reasoning tasks  (Huang and Chang, 2023). However, they still can not perform TR sufficiently well  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023). TR is more challenging since it requires a combination of various skills including mathematical and logical reasoning as well as commonsense knowledge  (Zhou et al., 2019; Qin et al., 2021).
Figure 1: Our TG-LLM framework has two steps: text-to-temporal graph translation and neuro-symbolic reasoning.
Recent works mainly adopt the general reasoning methods to investigate and improve the TR capability of LLMs. For example,  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023) benchmark leading LLMs on different TR tasks with standard prompting, few-shot learning and Chain of Thoughts (CoTs)  (Wei et al., 2023a).  (Li et al., 2023; Yuan et al., 2023) introduce pre-defined Python programs/rule-based templates to perform supervised fine-tuning (SFT). Similarly,  (Wei et al., 2023b) design several special types of prompts for in-context learning (ICL). In addition,  (Tan et al., 2023a, b) adopt some extra strategies, which include specific pre-training, instruction tuning and reinforcement learning.
Despite the effectiveness of such methods, they either ignore or not explicitly involve the intrinsic nature of TR. Human performs complex TR on a timeline of events which are aligned with the entities and relations. These temporal concepts (e.g., ordering, duration, frequency, typical time) are then rigorously defined based on the timeline information. Thus, a natural solution for LLMs to reach human-level performance is to first extract a timeline of events from the text, and then perform deliberate reasoning on it. However, the high-quality extraction of the aligned timeline is a challenging task itself due to the lack of ground truth. To solve this problem, we propose a synthetic dataset, which is fully controllable and requires minimal supervision. We show in experiments that the capability of timeline extraction learned on our dataset can be transferred to other tasks and benchmarks.
Given a reliable timeline, we still need to solve two problems to guide LLMs to perform rigorous reasoning on it: (1) How can one introduce the necessary arithmetic and commonsense knowledge involved in TR? (2) How can one guide LLMs to perform rigorous reasoning? Inspired by retrieval-augmented generation (RAG)  (Lewis et al., 2021), we first extract all the time expressions, and then generate all the necessary information (e.g., the temporal relation and time gap between the timestamps, the relative order of the gaps). As for (2), there are two roadmaps: (i) translating natural language into logical statements, and using external symbolic engine for reasoning  Pan et al. (2023); (ii) directly using LLMs as the reasoning engine  (Zhu et al., 2023). For (i), the difficulty lies in accurate translation  Yang et al. (2023c) and the limited expressive power of formal logic. For (ii), there is no guarantee for the correctness of the generated text especially when training data are insufficient  (Yang et al., 2023b). To solve these problems, we propose a symbolic-integration bootstrapping method to generate reliable intermediate steps for supervised fine-tuning. Furthermore, we design special data augmentation strategies to mitigate data deficiency for neuro-symbolic reasoning tasks.
To be specific, our contributions are summarized as follows:
  • We propose a new paradigm, TG-LLM, for language-based TR. In this framework, we first translate the context into a temporal knowledge graph, and then perform symbolic reasoning on it. Extensive experiments prove that our novel approach results in superior performance compared to the conventional approach.
  • We design two approaches including chain of thoughts bootstrapping and data augmentation to guide LLM to utilize symbolic reasoning, which brings better performance than free-form text.
  • We present a pipeline to create a synthetic dataset for question answering that requires TR. It is fully controllable and requires minimal supervision for text-temporal graph alignment. We show in experiments that fine-tuning on our dataset benefits LLMs on other TR tasks and benchmarks.
Temporal Graph (sub; rel; obj; start/end; time):
[…]
[2] (John Thompson owned Pearl Network) starts at 1942;
[3] (Sophia Parker was married to John Thompson) starts
at 1947; [4] (John Thompson was married to Sophia Parker)
starts at 1947; [5] (Sophia Parker was married to John
Thompson) ends at 1953; [6] (John Thompson was married
to Sophia Parker) ends at 1953; [7] (John Thompson owned
Pearl Network) ends at 1967; [8] (John Thompson died in
Riverside) starts at 1988; [9] (Sophia Parker died in
Lancaster) starts at 1995.
Graph-based Story (from GPT-3.5):
[…]
boy named John Thompson was brought into the world
in the year 1921. Growing up, he had a vibrant spirit
and an adventurous soul. $\cdots$
Graph-based QAs (from rule-based Python script):
[…]
Pearl Network) or (John Thompson was married to
Sophia Parker)?
A1: (John Thompson owned Pearl Network).
Q2: True or false: event (John Thompson owned Pearl
Network) was longer in duration than event (Sophia
Parker was married to John Thompson)?
A2: True.
$\cdots$
Table 1: Each sample from TGQA dataset is in the form of (story, temporal graph, questions, answers).
2 Dataset Construction
Reasoning Type Question Answer
Sequencing Given the following <N> events: <Event_A>, <Event_B>, <Event_C>, <Event_A>/<Event_B>/
<Event_D>, $\cdots$, which event is the first/second/third/fourth/$\cdots$ one in <Event_C>/<Event_D>/
the chronological order? $\cdots$
Duration How long did the event <Event_A> last? <Duration_of_Event_A>
Temporal How much time passed between the start of <Event_A> and the start of <Gap_between_Event_A
Relation <Event_B>? _and_B_start_time>
What happened right before/after <Event_A> started? <Event_B> right before/
after <Event_A>
Facts When did the <Event_A> occur? <Event_A_start_time>
Extraction
Simultaneity True or false: <Event_A> and <Event_B> happened at the same year? True / False
True or false: <Event_A> was still happening when <Event_B> started? True / False
Comparative True or false: <Event_A> was longer in duration than <Event_B>? True / False
Analysis
Table 2: Reasoning types with corresponding questions and answers in TGQA.
We propose a pipeline for the construction of our synthetic dataset called TGQA. Each sample, as shown in Table 1, is in the form of (story, temporal graph, questions, answers). The temporal graph (TG) is a subgraph extracted from a large temporal knowledge graph, YAGO11k (Dasgupta et al., 2018), with a restriction on the number of events. Each story is translated from the TG by GPT-3.5. Finally, by using some rule-based templates, we obtain reliable question and answer (QA) pairs from the graph. We obtain 500 samples for training/validation and another 100 for test, with about 30 QA pairs in a single sample. Compared with existing datasets  (Chen et al., 2021; Tan et al., 2023a), we have more diverse categories and types of temporal questions as shown in Table 2. Below we provide the details of each step in the dataset construction pipeline.
Step 1: Graph Splitting & Masking  Existing temporal knowledge graph datasets (Leetaru and Schrodt, 2013; Dasgupta et al., 2018; García-Durán et al., 2018) have tens of thousands of events, making it impossible to convert them into text all at once. Given a specific entity, we extract its third-order neighbors and all the events happening between them. Since we want LLMs to do reasoning instead of memorization, it is ensured that no overlapping exists between the facts in training/validation and test samples. In addition, we consider the noticeable corpus leakage problem of LLMs, where prior knowledge of test data has been implicitly obtained from the extensive pre-training process. Thus, a masking strategy, i.e., changing entity names into random ones of the same type, is adopted. We generate a global mapping for entity names with GPT-3.5 (see Table 6). To avoid confusion, we only choose obscure names that do not exist in the original large knowledge graph.
Step 2: Graph-based Open QA Creation  Based on the given TG, we generate a story and multiple QAs. We first ask GPT-3.5 to write a story based on the graph with the requirements to include all the facts. It is observed that GPT-3.5 tends to ignore the end time of some events in the created story. To solve this problem, we split the starting and ending of a persistent event into two transient ones (see Table 1). To obtain a comprehensive benchmark, we consider all types of temporal reasoning, which include sequencing, duration, frequency, simultaneity, temporal relation, comparative analysis and facts extraction. Given these categories, we design multiple diverse question types (see Table 2), and use a rule-based Python script to generate the corresponding Qs and As.
Step 3: Quality Control  Although LLMs are good at story generation, there is a risk of misalignment between the given TG and the generated text. To avoid the interference caused by misalignment, we propose a semi-automatic way for verification. Given a generated story, we ask LLMs the time of each event in the graph. If LLMs cannot give the correct answer, we consider the event as possibly missing in the story. Fully manual inspection is expensive, but by first asking LLM we can narrow down potential errors and only manually inspect the set of unanswered questions by LLM. Note that only the quality of story-TG alignment needs manual inspection, since all the QAs are generated from the TG with specific rules. We show dataset statistics in Appendix A, and all the prompts involved in Appendix C.
3 TG-LLM
Motivated by human perception, we propose a new paradigm called TG-LLM. We first translate the text into a temporal graph, and then guide the LLM to perform symbolic reasoning on it.
3.1 Text-to-TG Translation
TG serves as the foundation of the following symbolic reasoning process. Although LLMs can extract fact/events from the given context without fine-tuning, we observe a misalignment between the generated graph and pre-defined QAs, i.e, LLMs making mistakes or using different entity/relation names or time expressions. Thus, we provide a deliberate pipeline for high-quality TG translation.
Ground-truth TG Generation   For datasets such as TGQA, each story is corresponding to a verified TG. Thus, we can directly fine-tune LLMs on this task with the ground truth. However, the biggest challenge for most real-world applications is the lack of ground truth. Instead of using the naive estimation from LLMs, we provide a pipeline for high-quality generation.
(1) Entity & Relation Extraction: To align the entity/relation names in generated TGs and pre-defined QAs, we use a prompt with requirements on including the ones in pre-defined QAs when extracting from the context. To obtain the entity/relation names in pre-defined QAs, we consider two ways: parsing with rules or using LLMs.
(2) Temporal Info Identification: To identify all the time expressions, LLMs are first used to extract them from the context, and then a rule-based Python script is adopted for time normalization and invalid output filtering. For the valid normalized time expressions, we can easily obtain their temporal relations and time gaps, which are introduced as external knowledge in temporal reasoning (see Table 7).
(3) Graph Construction & Verification: We generate the graph using LLMs with ICL. To be specific, we provide in the prompt several examples which include a story and extracted entities, relations and time expressions as input and the corresponding TG as output (see Table 15). For those without explicit expressions, we ask LLMs to provide the temporal order of different events. We further semi-automatically verify the generated TGs by asking LLMs to answer the pre-defined QAs based on the graph. For those graphs where LLMs give the wrong answers, we manually check and correct them.
Figure 2: In chain of thoughts (CoTs) bootstrapping, we accept CoTs leading to correct answers and sample them according to their contrastive learning scores.
Supervised Fine-tuning  We found out in our experiments that providing a chronological list of events helps LLMs provide better temporal reasoning. In addition, as shown in Table 1, to deal with both transient and persistent events, we consider the starting and ending of a persistent event as two transient events. With these steps, we transfer the TG into a timeline with the alignment between entities, relations and times. We perform supervised fine-tuning (SFT) using Llama-2 model  (Touvron et al., 2023a) with Low-Rank Adaptation (LoRA)  (Hu et al., 2021). The input and output of the model are the story and aligned timeline, respectively. In experiments, we observe benefits from the SFT on our dataset to other temporal reasoning tasks.
3.2 Neuro-symbolic Temporal Reasoning
Given the reliable TGs, we train LLMs to perform symbolic reasoning with SFT enhanced by CoTs bootstrapping and data augmentation.
3.2.1 Bootstrapping Chain of Thoughts
It is observed that SFTs on reliable CoTs bring better performance in reasoning tasks than those on standard Input/Output prompts  (Wang et al., 2023; Ho et al., 2023). However, it is challenging to obtain reliable CoTs from LLMs for symbolic reasoning tasks  (Saparov and He, 2023). Inspired by  (Zhu et al., 2023), we provide several manually verified examples to LLMs as ICL, and ask LLMs to generate CoTs for the new samples.
To be specific, we first classify the pre-defined QAs into several categories $\{Q_{1},Q_{2},\cdots,Q_{M}\}$. For each category $Q_{i}$, we randomly choose $N_{i}$ samples $\{(g_{j},e_{j},q_{j},a_{j})\}^{N_{i}}_{j=1}$ and ask GPT-3.5/GPT-4 to provide symbolic CoTs $\{c_{j}\}$ given the TG $\{g_{j}\}$, external knowledge $\{e_{j}\}$ and question $\{q_{j}\}$. These CoTs will be manually verified and used as few-shot ICL examples. When given new samples $(g_{j^{\prime}},e_{j^{\prime}},q_{j^{\prime}},a_{j^{\prime}})$, we bootstrap CoTs with final answers $(c_{j^{\prime},k},\hat{a}_{j^{\prime},k})$ from GPT-3.5, and only accept the one leading to correct answers, i.e., $\hat{a}_{j^{\prime},k}=a^{*}_{j^{\prime}}$. In addition, inspired by  (Wang et al., 2023), we compare different CoTs using contrastive learning given in Eq. 2. The core idea is to choose CoTs that make the correct answer more plausible than other candidate answers. For the generated CoTs, we use weighted sampling according to the score-based probability given in Eq. 1.
$P_{sample}(c_{k})=\operatorname{softmax}(score(c_{k}))$ (1)
$score(c_{k})=\log P\left(a^{*}\mid g,e,q,c_{k}\right)+\gamma G\left(c_{k}\right)$ (2)
$G\left(c_{k}\right)=\log\frac{P\left(a^{*}\mid g,e,q,c_{k}\right)}{\bar{P}_{\{a^{\prime}\in A^{\prime}\}}\left(a^{\prime}\mid g,e,q,c_{k}\right)}$ (3)
$P\left(a\mid g,e,q,c_{k}\right)=\frac{1}{|a|}\underset{t_{l}\in a}{\sum}P\left(t_{l}\mid g,e,q,c_{k},t_{<l}\right)$ (4)
where $P_{sample}(\cdot)$ is the sampling probability for CoTs selection, $score(\cdot)$ is the contrastive learning score for each CoT, $g,e,q,c$ denote the TG, external knowledge, question, and CoT, respectively; $a^{*}$ denotes the correct answer; $a^{\prime}$ denotes the wrong answer from a candidate set $A^{\prime}$; weight $\gamma$ is a hyperparameter; $G(\cdot)$ denotes the contrastive plausibility growth; conditional probability $P(\cdot)$ is calculated in a way similar to perplexity, $t_{l}$ denotes the $l$-th token in the answer $a$, and $t_{<l}$ denotes the sequence of tokens before $t_{l}$.
Compared with generating CoTs with templates  (Saparov and He, 2023; Yuan et al., 2023), our method has more diversity and does not require prior knowledge for the reasoning process.
3.2.2 Data Augmentation
Compared with conventional tasks, reasoning suffers more from data insufficiency since more information, such as evidence, arguments, and logics, are involved in intermediate steps  (Huang and Chang, 2023). To address the issue, we propose special strategies for data augmentation.
Essentially, our framework involves two steps: text-to-TG translation and symbolic temporal reasoning. At the test stage, reasoning happens on the noisy estimated graphs, which impacts the overall performance. Specifically, for the reasoning module, we train with ground-truth/verified graphs while in the test time, the estimated graph is used for symbolic reasoning. To improve the robustness of reasoning, we introduce some disturbances on the graphs during training. Note that the type of disturbances should be carefully designed in order to avoid confusing the model. We first investigate two types of disturbances: (i) remove irrelevant edges and (ii) replace edges by changing relation names, given in Eq. 5 and Eq. 6, respectively. An edge is considered irrelevant if it is not involved in both QA and CoT.
$F_{irr}\in\{F:P(c|g_{\setminus\{F\}},e,q)\approx P_{0}\}_{F\in g}$ (5)
$F^{\prime}\in\{f_{R}(F):P(c|f_{R}(g),e,q)\approx P_{0}\}_{F\in g}$ (6)
Figure 3: We have several strategies for data augmentation: remove irrelevant edges, use relation synonyms and change entities/times.
where $P_{0}:=P(c|g,e,q)$ denotes the original conditional probability, irrelevant event $F_{irr}$ will be randomly removed from $g$, and disturbed event $F^{\prime}$ is obtained from the global mapping of relation names $f_{R}(\cdot)$.
Furthermore, to make sure LLM learns the underlying principle of symbolic temporal reasoning instead of just memorizing the training data, we introduce another two types of disturbances: (iii) globally map all the entity names in training data to some random names of the same type , and (iv) change the times based on a global offset, given in Eq. 7. For each relation, we generate these random names with GPT-3.5 by providing several examples of existing names.
$g^{\prime\prime}=f_{T}(f_{E}(g)),\ e^{\prime\prime}=f_{T}(e),$ (7)
$q^{\prime\prime}=f_{T}(f_{E}(q)),\ c^{\prime\prime}=f_{T}(f_{E}(c)),$
$a^{\prime\prime}=f_{T}(f_{E}(a))$
where $f_{E}(\cdot),f_{T}(\cdot)$ denote the global mapping of entity names and time changing, respectively.
4 Experiments
We aim to answer the following research questions in our experiments: (1) Can we guide LLMs to perform more reliable symbolic reasoning with the strategies of bootstrapping CoTs and data augmentation? (2) Can our two-step framework lead to better temporal reasoning performance? (3) Does these learned capabilities of temporal reasoning generalize to other tasks?
4.1 Experimental Setup
In addition to TGQA, we consider the two existing datasets, TimeQA  (Chen et al., 2021) and TempReason  (Tan et al., 2023a), which are constructed using Wikipedia articles, excerpts, or summaries. Compared with these datasets, TGQA has a better quality due to the controlled generation, and more diverse question categories (see examples from other datasets in Appendix B). We evaluate our framework on all the datasets with the metrics of token-level F1, exact match (EM) and perplexity-based accuracy (Acc). Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider Acc for LLM evaluations, i.e., selecting from a candidate set the answer with the lowest perplexity from LLMs as their prediction. (See rationales for evaluation metrics selection in Appendix B)
We compare our framework [^1] Code and data available at https://github.com/xiongsi heng/TG-LLM. with leading LLMs including Llama2 Touvron et al. (2023a), GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI et al., 2023). We run the Llama2 family on local machines (4-bit quantization for Llama2-70B due to limited computational resources), and use APIs provided by OpenAI for GPT models. We evaluate their few-shot ICL performance on the test set. To avoid LLMs generating free text which affects the evaluation, we provide candidate answers for all questions. The model versions and prompt templates are provided in Appendix B and C, respectively. We also include the T5-based models  (Tan et al., 2023a; Yang et al., 2023a) for a comprehensive comparison. Results on TimeQA and TempReason reported in the original papers are used, and these models are fine-tuned and evaluated on TGQA.
4.2 Implementation Details
We use Llama2-13B as the baseline due to limited computational resources. We inject two adapters with selectors into the model for the text-to-TG translation and symbolic temporal reasoning. The adapters are trained in parallel. For inference, we first translate the original story into a temporal graph, and then perform symbolic reasoning on it, i.e., the adapters are used in sequence. In addition, we use GPT-3.5/GPT-4 for all the generation of story, graph and CoTs due to their high quality generation. All the prompt templates are given in Appendix C.
Figure 4: Performance comparison between different strategies of CoT generation on TGQA.
Strategy ER-T1 ER-T2 ER-T3 ER-T4
ICL 0.13 0.13 0.31 0.10
SFT 0.04 0.17 0.13 0.10
SFT + bs 0.06 0.13 0.03 0.08
SFT + bs + aug 0.03 0.05 0.04 0.05
Table 3: Human evaluation on the CoTs generated by different strategies. T1: using wrong info; T2: logical inconsistency; T3: external knowledge error; T4: temporal graph error. (See Appendix B for error type explanations)
  Model TGQA TimeQA TempReason
Easy-mode Hard-mode OBQA-L2 OBQA-L3
EM F1 Acc EM F1 Acc EM F1 Acc EM F1 Acc EM F1 Acc
${}^{{\dagger}}$ 0.410 0.608 - 0.600 0.682 - 0.556 0.641 - 0.260 0.450 - 0.238 0.418 -
${}^{{\dagger}}$ 0.548 0.713 - 0.631 0.716 - 0.595 0.681 - 0.327 0.509 - 0.288 0.468 -
${}^{{\dagger}}$ 0.640 0.778 - - - - - - - 0.318 0.496 - 0.261 0.430 -
${}^{{\dagger}}$ 0.435 0.633 - 0.614 0.704 - 0.582 0.673 - 0.336 0.516 - 0.285 0.449 -
${}^{{\dagger}}$ 0.461 0.660 - 0.637 0.723 - 0.605 0.693 - 0.374 0.549 - 0.334 0.493 -
GPT-3.5 (ICL-SP) 0.598 0.699 - 0.640 0.668 - 0.512 0.506 - 0.303 0.409 - 0.365 0.453 -
GPT-3.5 (ICL-CoT) 0.706 0.788 - 0.565 0.603 - 0.436 0.464 - 0.340 0.478 - 0.243 0.348 -
$*{}^{*}$ (ICL-SP) 0.772 0.829 - 0.716 0.742 - 0.571 0.546 - 0.454 0.525 - 0.431 0.485 -
$*{}^{*}$ (ICL-CoT) 0.821 0.865 - 0.662 0.693 - 0.618 0.636 - 0.388 0.480 - 0.352 0.447 -
Llama2-7B (ICL-SP) 0.415 0.596 0.447 0.352 0.408 0.367 0.341 0.404 0.354 0.205 0.344 0.226 0.044 0.084 0.153
Llama2-7B (ICL-CoT) 0.548 0.686 0.578 0.367 0.425 0.393 0.302 0.354 0.360 0.233 0.410 0.263 0.179 0.357 0.187
Llama2-13B (ICL-SP) 0.440 0.609 0.526 0.439 0.493 0.450 0.427 0.481 0.437 0.284 0.452 0.289 0.189 0.370 0.183
Llama2-13B (ICL-CoT) 0.628 0.762 0.668 0.518 0.572 0.535 0.434 0.490 0.480 0.330 0.498 0.368 0.242 0.391 0.272
Llama2-70B (ICL-SP) 0.618 0.736 0.665 0.583 0.627 0.631 0.493 0.537 0.551 0.358 0.491 0.387 0.128 0.181 0.148
Llama2-70B (ICL-CoT) 0.761 0.838 0.796 0.552 0.612 0.623 0.447 0.501 0.512 0.325 0.477 0.345 0.303 0.393 0.318
Llama2-13B (SFT-SP) 0.550 0.720 0.652 0.412 0.449 0.455 0.362 0.404 0.412 0.337 0.506 0.338 0.244 0.408 0.253
Llama2-13B (SFT-CoT) 0.646 0.722 0.705 0.550 0.586 0.564 0.332 0.391 0.379 0.256 0.433 0.281 0.285 0.409 0.305
Llama2-13B (SFT-TGSR) 0.797 0.850 0.819 0.664 0.691 0.673 0.631 0.664 0.649 0.424 0.522 0.432 0.356 0.469 0.399
Table 4: Main results using different models and strategies. We report exact match (EM), token-level F1 scores, and perplexity-based accuracy (Acc). Note: (1) Results with ${}^{{\dagger}}$ are reported in the original papers. We fine-tune and evaluate the models on our dataset. (2) Results with * are evaluated on 1000 random test samples.
Figure 5: Performance comparison between different strategies on TimeQA and TempReason. To obtain a fair comparison, we use Llama2-13B as the base model for all strategies. The basic strategy is in-context learning with standard Input/Output (ICL-SP), which is used to calculate the performance changes.
4.3 Main Results
Can our strategy bring more reliable symbolic reasoning?   Figure 4 shows the comparison between ICL with CoTs, SFT with CoTs, bootstrapping CoTs, and data augmentation on TGQA. It can be seen that LLMs learn better temporal reasoning with SFT than ICL. By providing CoTs with bootstrapping and our data augmentation strategies, the model performance gets further enhanced. Furthermore, inspired by  (Wang et al., 2023), we manually check 100 CoTs generated by different strategies. Evaluators are asked to classify the errors into four types (using wrong info, logical inconsistency, external knowledge error, and temporal graph error). It can be seen in Table 3 that our strategies reduce all types of error rates (ER).
Can our two-step framework lead to better temporal reasoning performance?   Table 4 shows the comparison between different models on all datasets. First, we observed that among all the LLMs with ICL, GPT-4 has the strongest performance as expected. For the Llama2 family, larger models have better performance due to advanced context understanding and improved generalization. We also found that CoTs not always bring better temporal reasoning due to unreliable intermediate results and hallucinations. More importantly, our model (SFT-TGSR), which is based on Llama2-13B, shows a comparable or even better performance than GPT-4 on all datasets. To prove the effectiveness of our framework, we also provide the results of some alternative strategies, where SFT-SP and SFT-CoT denote supervised fine-tuning on standard Input/Output prompts and CoTs, respectively. It can be seen that the two-step framework brings a substantial improvement in model performance on different datasets. We hypothesize that this improvement in performance is because our two-step neuro-symbolic process provides an easier path toward answering the temporal questions for LLMs.
Does these learned capabilities of temporal reasoning generalize to other tasks?  Figure 5 shows the comparison between different strategies on TimeQA and TempReason. In our framework, LLMs can learn the capability of text-to-graph translation and symbolic reasoning, which brings better temporal reasoning. More importantly, we observe that per-training on TGQA improves the model performance compared with ICL. Also, it can be concluded that these necessary capabilities in temporal reasoning are generalizable to different data distributions. Since the synthetic dataset is fully controllable and requires minimal supervision, we actually provide a way of improving the temporal reasoning capability of LLMs.
4.4 Ablation Study
We ablate different modules to see their contributions to the performance. Table 5 shows the performance comparison between different configurations. To obtain a fair comparison, we use Llama2-13B as the base model for all configurations. From the ablation study, we obtain some insights: (1) LLMs can benefit from explicitly presented (temporal) graph which is intuitive, concise and structured. (2) Given a reliable graph, symbolic reasoning brings better performance than free-text CoTs. (3) Data augmentation is necessary for LLMs to perform complex tasks such as temporal reasoning. (4) The introduction of external knowledge such as mathematics and commonsense can augment the generation.
5 Related Work
Language-based Temporal Reasoning  Recently, language-based TR has gained substantial research focus  (Nylund et al., 2023; Hu et al., 2023; Gurnee and Tegmark, 2023; Liu et al., 2023). The vision here is to help LMs understand temporal concepts. Existing methods mainly solve this problem with time-aware modeling. For example,  (Rosin et al., 2022; Pereira, 2022; Tan et al., 2023a) propose pre-training/fine-tuning strategies for robust TR.  (Ning et al., 2019; Zhou et al., 2020; Yang et al., 2020; Zhou et al., 2021; Yang et al., 2023a) design auxiliary objectives to introduce external knowledge. Although these methods make some progress, the underlying structure and logic of TR are either ignored or not explicitly involved.
Reasoning towards LMs   Although LLMs exhibit some emergent behaviors, it is still unknown whether they can actually perform reasoning and how strong their capability of reasoning is (Huang and Chang, 2023). Existing methods that try to elicit or improve reasoning can be divided into two directions: reasoning-involved modeling or hybrid methods. For example,  (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023) use prompt engineering where some examples including intermediate reasoning steps are provided in context.  (Talmor et al., 2020; Nye et al., 2021; Wang et al., 2023; Ho et al., 2023) fine-tune LMs with these information before producing final answers. On the other hand,  (Li et al., 2023; Kynoch et al., 2023; Zhuang et al., 2023; Xu et al., 2023) combine LMs with domain-specific tools, empowering the model to perform more complex tasks that require reasoning.
Neuro-symbolic Reasoning Recent research has shown the advantages of combining neural networks with symbolic reasoning  (Sarker et al., 2021; Hamilton et al., 2022). In particular, knowledge graphs (KGs) as the foundation of (temporal) symbolic reasoning has been widely adopted in the past (Yang et al., 2017; Payani and Fekri, 2019a, b; Yang and Song, 2020; Qu et al., 2021). Moreover,  (Liu et al., 2022; Xiong et al., 2022, 2023) develop approaches for symbolic reasoning over temporal KGs. In this paper, we build the connection between language-based and temporal KG-based temporal symbolic reasoning. This connection brings the potential for extending these KG-based methods to language-based tasks. Different from existing methods (Luo et al., 2023; Zhang et al., 2023; Yuan et al., 2023), which limit their application to certain datasets, our framework offers enhanced generalization and usability, largely attributed to its innovative use of text-to-graph translation as a precursor to symbolic reasoning.
Config. EM F1 Acc
Baseline (SFT-CoT) 0.646 0.722 0.705
TG 0.675 0.754 0.711
TG + SR(bs) 0.723 0.797 0.736
TG + SR(bs + aug) 0.782 0.838 0.795
TG + EK + SR(bs) 0.753 0.803 0.776
TG + EK + SR(bs + aug) 0.797 0.850 0.819
Table 5: Ablation study results on TGQA. To obtain a fair comparison, we use Llama2-13B as the base model for all configurations. TG: temporal graph, SR: symbolic reasoning, EK: external knwoledge, bs: bootstrapping, aug: data augmentation.
6 Conclusion
TG-LLM, a new framework for language models, has been proposed to improve their performance on temporal reasoning. Besides producing the final answers, our framework also provides the temporal graph and symbolic reasoning steps. Experiments indicate that TG-LLM achieves better performance than existing methods. An interesting direction for future work is to extend it to more complex applications such as induction and abductive reasoning. Due to the graph structure and capability of symbolic reasoning, it is promising to improve the performance of LLMs on these tasks.
Limitations
Graph-augmented approaches such as TG-LLM help language models better learn related concepts from the perspective of graph. Although we demonstrate TG-LLM has good performance on understanding temporal relations, it still needs adaptations for temporal commonsense reasoning  (Zhou et al., 2019; Qin et al., 2021). Explicit in-context integration of commonsense presents opportunities for this task. Further, we mainly improve the capability of LLMs by introducing a new paradigm and providing more plausible and informative training data. There can be opportunities such as simulating an environment to provide feedback to LLMs  (Yao et al., 2023). For example, we can verify the generated TG based on prior knowledge such as the time gap between someone’s birth date and death date, i.e., a person’s lifespan, should fall into a certain range. In this way, we can further improve the quality of generation.
Ethics Statement
In this paper, we adopt YAGO11k for fine-tuning the language models. The dataset is publicly available, and is for research purposes only. We also use GPT model to generate text based on YAGO11k, for which OpenAI has been committed to addressing ethical considerations. In addition, we adopt TimeQA and TempReason for evaluation. Both datasets are publicly available, and are for research purposes only. However, they may still contain improper or harmful content. None of such content reflects the opinions of the authors.
Appendix A TGQA Dataset Statistics
In this section, we show statistics for our TGQA dataset. Figure 6 shows the number of facts distribution in training and test set, which determine the complexity of the task to some extent. Figure 7 shows the types of questions distribution in training and test set, where Q0: "Which event occurred first, <Event_A> or <Event_B>?", Q1: "Given the following <N> events: <Event_A>, <Event_B>, <Event_C>, <Event_D>, $\cdots$, which event is the first/second/third/fourth/$\cdots$ one in the chronological order?", Q2: "How long did the event <Event_A> last?", Q3: "", Q4: "How much time passed between the start of <Event_A> and the start of <Event_B>?", Q5: "What happened right before/after <Event_A> started?", Q6: "When did the <Event_A> occur?", Q7: "True or false: <Event_A> and <Event_B> happened at the same year?", Q8: "True or false: <Event_A> was still happening when <Event_B> started?". Note that we give eight categories of question in Table 2, since in a strict sense Q0 and Q1 can be considered as the same type. Among all the question types, Q1 has the largest portion since it has multiple variants. Similarly, Q7 with two variants also has a larger portion. To mitigate question category imbalance, we first calculate the metrics for each category, and then use the average as final scores. We also provide examples for the global mapping for entity names (Table 6) and external knowledge (Table 7) used in our dataset.
Entity Name Mapping
Ori New
Beverly_Adams Isabella_Thompson
Edmonton Lancaster
Al_Gore Chris_Evans
Carrara_Stadium Maplewood_Arena
Bend_Sinister_(novel) Lakefield_Chronicles_(novel)
$\cdots$ $\cdots$
Table 6: We use a global mapping for entity names from GPT-3.5 to avoid corpus leakage.
External Knowledge
1885 before 1893 before 1916 before 1918 before 1922
before 1928 before 1941
1918 - 1916 = 2
1928 - 1893 = 35
1928 - 1922 = 6
1941 - 1918 = 23
2 < 6 < 23 < 35
Table 7: We integrate the necessary mathematics in context as external knowledge for temporal reasoning.
Appendix B Experiment Details
In this section, we present more experiment details for further research. We include data augmentation examples, model versions and evaluation tasks & metrics.
Temporal Graph:
[…]
[1] (John Thompson was born in Weston) starts at 1921;
[2] (John Thompson run Pearl Network) starts at 1942;
[3] (Sophia Parker and John Thompson became life
partner) starts at 1947; [4] (John Thompson and Sophia
Parker became life partner) starts at 1947; [5] (Sophia
Parker and John Thompson became life partner)
ends at 1953; [6] (John Thompson and Sophia Parker)
became life partner) ends at 1953; [7] (John Thompson
run Pearl Network) ends at 1967.
Graph-based QAs:
[…]
Q: True or false: event (John Thompson owned Pearl
Network) was longer in duration than event (Sophia
Parker was married to John Thompson)?
CoT (symbolic): The duration for each event can be
calculated as follows:
(John Thompson owned Pearl Network) starts at
1942, ends at 1967, 1967 - 1942 = 25
(Sophia Parker was married to John Thompson) starts
at 1947, ends at 1953, 1953 - 1947 = 6
25 is greater than 6 , thus, the answer is True.
A: True.
Temporal Graph:
[…]
[1] (James Brown was born in Oslo) starts at 1931;
[2] (James Brown owned Iris Inn) starts at 1952;
[3] (Ella Perry was married to James Brown) starts
at 1957; [4] (James Brown was married to Ella Perry)
starts at 1957; [5] (Ella Perry was married to James
Brown) ends at 1963; [6] (James Brown was married
to Ella Perry) ends at 1963; [7] (James Brown owned
Iris Inn) ends at 1977; [8] (James Brown died in
Auckland) starts at 1998; [9] (Ella Perry died in
Monaco) starts at 2005.
Graph-based QAs:
[…]
Q: True or false: event (James Brown owned Iris Inn)
was longer in duration than event (Ella Perry was
married to James Brown)?
CoT (symbolic): The duration for each event can be
calculated as follows:
(James Brown owned Iris Inn) starts at 1952, ends at
1977, 1977 - 1952 = 25
(Ella Perry was married to James Brown) starts at
1957 , ends at 1963, 1963 - 1957 = 6
25 is greater than 6 , thus, the answer is True.
A: True.
Table 8: We propose special data augmentation strategies for symbolic reasoning over (temporal) graphs.
Data Augmentation Examples  We show with an example our data augmentation strategies in Table 8. For TGs with relation synonym replacement or irrelevant edges removal, there is no need for change on Graph-based QAs. For TGs with entity names/times mapping, we also need to change with the corresponding entity names/times in Graph-based QAs to ensure the consistency.
Model Versions  All the model versions of LLMs used in our experiments are listed below:
  • Llama2-7B (meta-llama/Llama-2-7b-hf)
  • Llama2-13B (meta-llama/Llama-2-13b-hf)
  • Llama2-70B (meta-llama/Llama-2-70b-hf)
  • GPT-3.5 (gpt-3.5-turbo)
  • GPT-4 (gpt-4-1106-preview)
For the Llama2 family, all the model weights can be downloaded from the platform of Hugging Face. For the GPT models, all the model weights can be accessed through the OpenAI APIs.
Evaluation Tasks & Metrics  In addition to TGQA, we consider the existing datasets TimeQA  (Chen et al., 2021) and TempReason  (Tan et al., 2023a), which are constructed using Wikipedia articles, excerpts, or summaries. To be specific, TimeQA  (Chen et al., 2021) contains two difficulty levels (see Table 9). The easy-level split tends to be information extraction while the hard-level split involves understanding the relation between different temporal expressions. TempReason  (Tan et al., 2023a) contains three levels (L1: Time-Time Relation, L2: Time-Event Relation, L3: Event-Event Relation) and three settings (OBQA: open-book QA, CBQA: close-book QA, ReasonQA: facts-based QA) (see Table 10). We observe that some stories in TempReason are incomplete which partially leads to the low accuracy of LLMs. We evaluate our framework on all the datasets with the metrics of token-level F1, exact match (EM) and perplexity-based accuracy (Acc). Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider Acc for LLM evaluations.
TimeQA
George Washington (February 22, 1732 – December 14,
1799) was an American Founding Father, military offic-
er, politician and statesman who served as the first $\cdots$
Questions (Easy-mode):
What position did George Washington hold in June 17-
75?
What position was held by George Washington between
1778 and 1788?
$\cdots$
Questions (Hard-mode):
George Washington took which position before 1778?
What was George Washington’s position in early 1780s?
$\cdots$
Table 9: Example questions of two difficulty levels in TimeQA. Easy-mode: the query time expression is explicitly mentioned in the story. Hard-mode: obtaining the answer needs inference based on the temporal relation between the query time expression and the one mentioned in the story.
TempReason
Lionel Andrés "Leo" Messi (born 24 June 1987) is an Ar-
gentine professional footballer who plays as a forward for
and captains both Major League Soccer club $\cdots$
Questions (L1):
What is the year after 2010?
$\cdots$
Questions (L2):
What team did Leo Messi play for in 2010?
$\cdots$
Questions (L3):
What team did Leo Messi play for after Barcelona?
$\cdots$
Table 10: Example questions of three difficulty levels under the OBQA setting in TempReason. L1: Time-Time Relation, L2: Time-Event Relation, L3: Event-Event Relation.
Compared with other LMs, LLMs are pre-trained on a wide variety of tasks, making them more generalizable at the cost of less conforming to the required format. This property actually hurts their performance under the commonly used metrics such as EM and F1. To solve this problem, we select from a candidate set the answer with the lowest perplexity from LLMs as their prediction. For questions with multiple correct answers, we follow the strategy proposed in  (Chen et al., 2021) to get the best F1, EM and Acc among them. Since there are multiple categories of questions in TGQA, we first calculate the metrics for each category, and then use the average as final scores to mitigate question category imbalance.
  Model TGQA
Q0 Q1 Q2 Q3 Q4 Q5 Q6 Q7 Q8 Total
${}^{\ddagger}$ (0-shot SP) 0.801 0.435 0.660 0.355 0.930 0.955 0.710 0.315 0.617 0.642
GPT-3.5 (1-shot SP) 0.842 0.411 0.656 0.555 0.924 0.650 0.500 0.353 0.495 0.598
${}^{\ddagger}$ (0-shot CoT) 0.930 0.565 0.680 0.665 0.920 0.945 0.695 0.395 0.596 0.710
GPT-3.5 (1-shot CoT) 0.884 0.457 0.783 0.791 0.921 0.832 0.755 0.357 0.574 0.706
${}^{{\ddagger}*}$ (0-shot SP) 0.801 0.689 0.920 0.620 0.930 0.940 0.758 0.575 0.737 0.776
$*{}^{*}$ (1-shot SP) 0.820 0.660 0.870 0.630 0.910 0.990 0.760 0.610 0.700 0.772
${}^{{\ddagger}*}$ (0-shot CoT) 0.950 0.692 0.901 0.881 0.919 0.980 0.778 0.647 0.858 0.848
$*{}^{*}$ (1-shot CoT) 0.930 0.758 0.903 0.869 0.892 0.920 0.758 0.667 0.680 0.821
Llama2-7B (1-shot SP) 0.554 0.216 0.163 0.147 0.353 0.924 0.683 0.115 0.579 0.415
Llama2-7B (1-shot CoT) 0.660 0.264 0.484 0.603 0.693 0.947 0.632 0.079 0.574 0.548
Llama2-13B (1-shot SP) 0.442 0.341 0.353 0.192 0.686 0.640 0.561 0.163 0.580 0.440
Llama2-13B (1-shot CoT) 0.597 0.263 0.670 0.678 0.881 0.947 0.736 0.258 0.622 0.628
Llama2-70B (1-shot SP) 0.752 0.447 0.520 0.349 0.941 0.891 0.665 0.413 0.585 0.618
Llama2-70B (1-shot CoT) 0.980 0.490 0.878 0.832 0.944 0.878 0.807 0.294 0.750 0.761
Llama2-13B (1-shot SFT - TG) 0.878 0.515 0.656 0.866 0.835 0.845 0.726 0.369 0.383 0.675
Llama2-13B (1-shot SFT - TG + SR(bs)) 0.947 0.732 0.747 0.825 0.931 0.987 0.495 0.466 0.649 0.753
Llama2-13B (1-shot SFT - TG + SR(bs + aug)) 0.931 0.710 0.729 0.860 0.927 0.977 0.627 0.534 0.739 0.782
Llama2-13B (1-shot SFT - TG + EK + SR(bs + aug)) 0.944 0.775 0.729 0.849 0.924 0.980 0.783 0.506 0.681 0.797
[…]
We explain the categories of errors in human evaluation on CoT generation here (see Table 3). T1 means that LLM uses wrong information such as wrong start/end time during reasoning. T2 means that LLM makes logical errors, e.g., after determining "Event A ends before Event B starts", considering the statement as true that "Event A was still happening when Event B starts". T3 means that LLM makes errors on external knowledge, e.g., claiming that "the date 1978 is after 1983". T4 means that there exists error in the extracted TG which leads to the wrong conclusion.
Appendix C Example Prompts
In this section, we show example prompts used in our framework. To be specific, Table 12 shows an example of the graph-based story generation in TGQA; Table 13 presents an example of the automatic story-temporal graph alignment verification in TGQA; Table 14 provides an example of the temporal info identification in TimeQA; Table 15 illustrates an example of the graph construction in TimeQA; Table 16 gives an example of the automatic temporal graph-QA alignment verification in TimeQA; Table 17 depicts an illustrative case of the CoTs bootstrapping in TGQA.
Appendix D Fine-grained Results
We provide fine-grained results on TGQA in Table 11. In TGQA, there are nine types of questions as explained in Appendix A. We show all the models with different strategies here. Zero-shot performance of LLMs is also considered to investigate the effect of in-context examples. Note that we do not include zero-shot performance of the Llama2 family since it is not fine-tuned on instructions, i.e., we can not obtain valid zero-shot learning results. For zero-shot learning results, the format of generation is not guaranteed. The original accuracy is very low since rule-based parser cannot handle it. Instead, we use GPT-3.5 as parser to extract the answer from the generation. Also, we only adopt one example in context due to context length limitation.
It can be seen that in-context examples are not necessary for powerful pre-trained models such as GPT-3.5 and GPT-4. Undoubtedly, GPT-4 obtains the best in-context learning performance on most categories. However, for some categories, the best performance of GPT-4 is around 0.6 or 0.7, which is not sufficiently well. On the other hand, we show that our strategies improve the performance step by step. Note that our model achieves the best overall performance by improving all the categories. To be specific, symbolic reasoning brings better performance for most categories. Furthermore, data augmentation and external knowledge further enhance the capability of the model while preserving the existing good performance on those categories.
(Molly Adams was born in Seattle) starts at 1896
(Liam Thomas Dawson was born in Seattle) starts at 1896
(Molly Adams was married to Liam Thomas Dawson) starts at 1920
(Liam Thomas Dawson was married to Molly Adams) starts at 1920
(Liam Thomas Dawson won prize James Parker Prize for Advanced Biomedical
Studies) starts at 1946
(Liam Thomas Dawson won prize Champion Award in Biology or Science) starts at 1947
(Liam Thomas Dawson won prize Oakley Smith Prize) starts at 1948
(Liam Thomas Dawson won prize Member of the National Academy) starts at 1950
(Molly Adams was married to Liam Thomas Dawson) ends at 1957
(Liam Thomas Dawson was married to Molly Adams) ends at 1957
(Liam Thomas Dawson died in Kingsville, Texas) starts at 1984
Based on the above facts, generate a short story for me.
Table 12: Example of the graph-based story generation in TGQA.
Once upon a time, in the vibrant city of Seattle, two remarkable individuals, Molly
Adams and Liam Thomas Dawson, were born in the year 1896. Little did they know that
their destinies were entwined from the very beginning.
Growing up in the same neighborhood, Molly and Liam developed a deep friendship
that blossomed into something more as they entered adulthood. In the year 1920, their
love story officially began as they exchanged vows and embarked on a journey of
companionship that would last for decades.
$\cdots$
When did the event (Molly Adams was married to Liam Thomas Dawson) end?
Table 13: Example of the automatic story-temporal graph alignment verification in TGQA.
Knox Cunningham
Sir Samuel Knox Cunningham, 1st Baronet, QC (3 April 1909 – 29 July 1976) was a
Northern Irish barrister, businessman and politician. As an Ulster Unionist politician at
a time when the Unionists were part of the Conservative Party, he was also a significant
figure in United Kingdom politics as Parliamentary Private Secretary to Harold
Macmillan. His nephew was Sir Josias Cunningham.
$\cdots$
Extract all the time expressions such as ’June 1994’, ’1973’, ’late 1980s’.
Table 14: Example of the temporal info identification in TimeQA.
Knox Cunningham
Sir Samuel Knox Cunningham, 1st Baronet, QC (3 April 1909 – 29 July 1976) was a
Northern Irish barrister, businessman and politician. As an Ulster Unionist politician at
a time when the Unionists were part of the Conservative Party, he was also a significant
figure in United Kingdom politics as Parliamentary Private Secretary to Harold
Macmillan. His nephew was Sir Josias Cunningham.
$\cdots$
Construct a timeline for Knox Cunningham’s position. You should only consider these
time points (3 April 1909, 1930s, 1939, 1942, 1943, 1945, 1947, 1949, 1954, 29 July 1976).
Table 15: Example of the graph construction in TimeQA.
3 April 1909: Knox Cunningham was born.
1930s: He studied law and began his legal career.
1935: On 2 July 1935, he married Dorothy Enid Riley JP.
1939: Cunningham was called to the Bar by the Middle Temple.
1942: He was called to the Bar in Northern Ireland.
1943: Cunningham fought the Belfast West by-election.
1945: He contested the same seat in the general election.
1947: Knox Cunningham became involved with the World Alliance of YMCAs.
1949: He became Chairman of the National Council of the YMCA.
1954: Cunningham was elected to Orpington Urban District Council.
29 July 1976: Knox Cunningham passed away.
Which position did Knox Cunningham hold from May 1955 to Apr 1956?
Table 16: Example of the automatic temporal graph-QA alignment verification in TimeQA.
Example:
Timeline:
(Liam Mitchell was born in Harrison) starts at 1885
(Maddox Reynolds was born in Glenville) starts at 1893
(Liam Mitchell was married to Emma Scott) starts at 1916
(Emma Scott was married to Liam Mitchell) starts at 1916
(Liam Mitchell was married to Emma Scott) ends at 1918
(Emma Scott was married to Liam Mitchell) ends at 1918
(Maddox Reynolds was married to Liam Mitchell) starts at 1922
(Liam Mitchell was married to Maddox Reynolds) starts at 1922
(Maddox Reynolds was married to Liam Mitchell) ends at 1928
(Liam Mitchell was married to Maddox Reynolds) ends at 1928
(Liam Mitchell died in Boston) starts at 1941
(Emma Scott died in Oceanview) starts at 1984
Question: True or false: event (Liam Mitchell was married to Maddox Reynolds) was
longer in duration than event (Emma Scott was married to Liam Mitchell)?
Useful information:
1885 before 1893 before 1916 before 1918 before 1922 before 1928 before 1941
1918 - 1916 = 2
1928 - 1893 = 35
1928 - 1922 = 6
1941 - 1918 = 23
2 < 6 < 23 < 35
Answer: Let’s think step by step.
To determine whether the statement is true or false, we need to compare the durations
of (Liam Mitchell was married to Maddox Reynolds) and (Emma Scott was married to
Liam Mitchell) . The duration for each event can be calculated as follows:
(Liam Mitchell was married to Maddox Reynolds) :
(Liam Mitchell was married to Maddox Reynolds) starts at 1922
(Liam Mitchell was married to Maddox Reynolds) ends at 1928
1928 - 1922 = 6
(Emma Scott was married to Liam Mitchell) :
(Emma Scott was married to Liam Mitchell) starts at 1916
(Emma Scott was married to Liam Mitchell) ends at 1918
1918 - 1916 = 2
Now, we compare the duration of the two events: 6 is greater than 2 , thus, the answer is True .
Test:
Timeline:
(Chris Evans was born in Bristol, Connecticut) starts at 1948
(Oliver Jensen was born in Bristol, Connecticut) starts at 1948
(Chris Evans was married to Oliver Jensen) starts at 1970
(Oliver Jensen was married to Chris Evans) starts at 1970
(Chris Evans created Pearl Network) starts at 2005
(Chris Evans won prize Victory Achievement Award) starts at 2007
(Chris Evans was married to Oliver Jensen) ends at 2010
(Oliver Jensen was married to Chris Evans) ends at 2010
(Chris Evans created Pearl Network) ends at 2013
Question: True or false: event (Chris Evans owned Pearl Network) was longer in
duration than event (Oliver Jensen was married to Chris Evans)?
Useful information:
1948 before 1970 before 2005 before 2007 before 2010
1970 - 1948 = 22
2007 - 1948 = 59
2007 - 2005 = 2
2010 - 1970 = 40
2 < 22 < 40 < 59
Answer: Let’s think step by step.
Table 17: Example of the CoTs bootstrapping in TGQA.